I need some help with log please
Solve ln(3+3x)-ln(x-4)=ln5
thanks for your time

Answer by uma(370)

ln(3+3x) - ln(x-4) = ln5
==> ln[(3+3x)/(x-4)] = ln 5 [because lna - ln b = ln(a/b)]
Removing ln from either side we get....
[(3+3x)/(x-4)] = 5
Multiplying by (x-4) on either side we get..
3 + 3x = 5(x-4)
==> 3 + 3x = 5x - 20
==> 3 + 3x - 5x = 5x - 20 - 5x
==> 3 - 2x = - 20
==> 3 - 2x - 3 = - 20 - 3
==> -2x = - 23
==> -2x/-2 = -23/-2
==> x = 23/2
==> x = 11.5

Good Luck!!!